A laboratory believes that its equipment needs adjustment when the standard deviation of five samples of a known concertationWhich test do we use and why?
Answer #1

Solution:

Solution:

Confidence interval for population standard deviation is given as below:

Sqrt[(n – 1)*S2 / χ2α/2, n– 1 ] < σ < sqrt[(n – 1)*S2 / χ21 -α/2, n– 1 ]

We assume Confidence level = 95%

From given data, we have

Sample size = n = 5

Degrees of freedom = n – 1 = 4

Sample standard deviation = S = 4.8844

χ2α/2, n – 1 = 11.1433

χ21 -α/2, n– 1 = 0.4844

(By using chi square table)

Sqrt[(5 – 1)* 4.8844^2/ 11.1433] < σ < sqrt[(5 – 1)* 4.8844^2/ 0.4844]

Sqrt[8.5639] < σ < sqrt[196.9979]

2.9264 < σ < 14.0356

Lower limit = 2.9264

Upper limit = 14.0356

Confidence interval does not suggest that the equipment requires adjustment, because the above interval contains the value 5.5.